# A “Staircase” formula for the Chern-Schwartz-MacPherson cycle of a matroid

### v2

@@ 4 Relation to the CSM cycle of a matroid
Given a loopless matroid $M$, define the invariant $\operatorname{stair}(M)$ in its Chow ring $A^{*}(M)$ by the formula
@@
First we show that this invariant respects certain deletion and contraction operations on a matroid.
@@
Let $M$ be a matroid on ground set $E$ with rank function $\operatorname{rk}_{M}$. For any $i\in E$, we can construct the matroid $M\setminus i$, called the deletion of $M$ by the element $i$. This is the matroid on the set $E\setminus i$ with rank function $\operatorname{rk}_{M\setminus i}:2^{E\setminus i}$
@@
For any $i\in E$, we can construct the matroid $M/i$, called the contraction of $M$ by the element $i$. This is the matroid on the set $E\setminus i$ with rank function $\operatorname{rk}_{M\setminus i}:2^{E\setminus i}$
@@ 4.1 Contraction-deletion of $\operatorname{stair}$
For a given element $i\in E$ which is not a coloop, we obtain a projection $\delta:B(M)\to B(M\setminus i)$. By Proposition 2.22 in [Sha13], $\delta$ is a tropical modification along some rational function $\phi$; therefore, $\delta$ induces a pullback $\delta^{*}:Z_{*}(B(M\setminus i)\to Z_{*}(B(M))$ and since $div_{B\setminus i}(\phi)=B(M/i)$, we also obtain a pullback $Z_{*}(B(M/i))\to Z_{*}(B(M\setminus i))$. These have corresponding ring/group homomorphisms in the context of the Chow rings, see Section 4.2 for details.
@@
To distinguish the divisors of $A^{*}(M)$, $A^{*}(M\setminus i)$, and $A^{*}(M/i)$, we label them by
@@
$A^{*}(M/i)$ $=\mathbb{Z}[z_{F}\mid F\in\mathcal{L}(M/i)]/(I_{M/i}+J_{M/i})$
i.e. with variable name $x$, $y$, $z$ respectively. Define the deletion pullback (a ring homomorphism) $\bar{\delta}^{*}:A^{*}(M\setminus i)\rightarrow A^{*}(M)$ on each generator $y_{F}$ by
@@
where $\iota(z_{F})=y_{F}$ (which replaces the variables $z$ with $y$). This is an embedding $A^{*}(M/i)$ as the subgroup $y_{i}\cdot A^{*}(M\setminus i)$:
By composition we also obtain a map $\bar{\delta}^{*}(y_{i}\cdot\iota(\_)):A^{*}(M/i)\rightarrow A^{*}(M)$.
@@
The invariant $\operatorname{stair}(M)$ satisfies the contraction-deletion formula:
$\operatorname{stair}(M)=\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))-\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/i)))$
Proof.
We partition the nonempty flats of $M$ according to the flowchart:
[…]
Denote also $\mathcal{U}\setminus i:=\{F\setminus i|F\in\mathcal{U}\}$ and $\mathcal{V}\setminus i:=\{F\setminus i|F\in\mathcal{V}\}$.
Note the following reasons for this partition:
  • $\mathcal{U}\cup\mathcal{V}$ consists of all flats containing the deleted element $i$. Among these, $\mathcal{U}$ are the flats that decrease in rank when $i$ is deleted. On the other hand, $\mathcal{V}$ are the flats that remain the same rank when $i$ is deleted.
  • $\mathcal{U}\setminus i=\mathcal{S}\cup\{\emptyset\}$.
  • If $F\in\mathcal{T}$, then $\bar{\delta}^{*}(y_{F})=x_{F}$.
  • If $F\in\mathcal{S}$, then $\bar{\delta}^{*}(y_{F})=x_{F}+x_{F\cup\{i\}}$.
  • If $F\in\mathcal{V}\setminus i$, then $\bar{\delta}^{*}(y_{F})=x_{F\cup i}$.
[…]
For ranks $r=1,\dots,d+1$, let
The one piece may be empty in some rank $r$, in which case the sum is zero. For short hand, we also group indices with the same rank $(t+s)_{r}:=t_{r}+s_{r}$, etc.
Using this notation,
$\operatorname{stair}(M)$ $=[1-(t+s+u+v)_{1}-(t+s+u+v)_{2}-\dots-(t+s+u+v)_{d+1}]$
$\quad\quad\cdot[1-(t+s+u+v)_{2}-\dots-(t+s+u+v)_{d+1}]$
[…]
$\quad\quad\cdot[1-(t+s+u+v)_{d+1}]$
and
$\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))$ $=[1-(t_{1}+(s_{1}+u_{2})+v_{1})-(t_{2}+(s_{2}+u_{3})+v_{2})-\dots-(t_{d+1}+(s_{d+1}+u_{d+2})+v_{d+1})]$
$\quad\quad\cdot[1-(t_{2}+(s_{2}+u_{3})+v_{2})-(t_{3}+(s_{3}+u_{4})+v_{3})-\dots-(t_{d+1}+(s_{d+1}+u_{d+2})+v_{d+1})]$
[…]
$\quad\quad\cdot[1-(t_{d+1}+(s_{d+1}+u_{d+2})+v_{d+1})]$
$=[1-(t+s+v)_{1}-(t+s+u+v)_{2}-\dots-(t+s+u+v)_{d+1}]$
$\quad\quad\cdot[1-(t+s+v)_{2}-(t+s+u+v)_{3}-\dots-(t+s+u+v)_{d+1}]$
[…]
$\quad\quad\cdot[1-(t+s+v)_{d+1}]$
Note $u_{r}$ is omitted from the $i$-th factor. We compute the difference $\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))$ and show it is equal to the negative of
$\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/i)))=$
$[u_{1}+u_{2}+\dots+u_{d+1}]$
$\quad\cdot[1-(s_{1}+u_{2}+v_{2})-(s_{2}+u_{3}+v_{3})-\dots-(s_{d}+u_{d+1}+v_{d+1})]$
$\quad\cdot[1-(s_{2}+u_{3}+v_{3})-\dots-(s_{d}+u_{d+1}+v_{d+1})]$
[…]
$\quad\cdot[1-(s_{d}+v_{d+1}+u_{d+1})]$
The difference is:
$[-u_{1}]\cdot[1-(t+s+u+v)_{2}-\dots-(t+s+u+v)_{d+1}]\cdots[1-(t+s+u+v)_{d+1}]$
$+[1-(t+s+v)_{1}-(t+s+u+v)_{2}-\dots-(t+s+u+v)_{d+1}]\cdot[-u_{2}]\cdots[1-(t+s+u+v)_{d+1}]$
+$\cdots$+ ⋯
$+[1-(t+s+v)_{1}-\dots-(t+s+u+v)_{d+1}]\cdot[1-(t+s+v)_{2}-\dots-(t+s+u+v)_{d+1}]\cdots[-u_{d+1}]$
Note in the $r$-th term is like $\operatorname{stair}(M)$ except the $r$-th factor is replaced by $-u_{r}$ and for each $k<r$, $u_{k}$ each omitted from the $k$-th factor. Each flat in $\mathcal{T}$ is not comparable with each flat in $\mathcal{U}$, so $t_{k}u_{r}=0$. In other words, $t_{k}$ is annihilated by each $u_{r}$. Since each term has some $u_{r}$ as a factor, all ‘$t$’s may be removed.
$\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))=\dots=$
$[-u_{1}]\cdot[1-(s+u+v)_{2}-\dots-(s+u+v)_{d+1}]\cdots[1-(s+u+v)_{d+1}]$
$+[1-(s+v)_{1}-(s+u+v)_{2}-\dots-(s+u+v)_{d+1}]\cdot[-u_{2}]\cdots[1-(s+u+v)_{d+1}]$
+$\cdots$+ ⋯
$+[1-(s+v)_{1}-\dots-(s+u+v)_{d+1}]\cdot[1-(s+v)_{2}-\dots-(s+u+v)_{d+1}]\cdots[-u_{d+1}]$
Also $u_{r}s_{k}=0$ if $k\geq r$. Then, for each term, index $r$, for all $k>r$ add $s_{k-1}$ to the $k$-th factor (after each $[-u_{r}]$).
$\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))=\dots=$
$[-u_{1}]\cdot[1-s_{1}-(s+u+v)_{2}-\dots-(s+u+v)_{d+1}]\cdots[1-s_{d}-(s+u+v)_{d+1}]$
$+[1-(s+v)_{1}-(s+u+v)_{2}-\dots-(s+u+v)_{d+1}]\cdot[-u_{2}]\cdots[1-s_{d}-(s+u+v)_{d+1}]$
+$\cdots$+ ⋯
$+[1-(s+v)_{1}-\dots-(s+u+v)_{d+1}]\cdot[1-(s+v)_{2}-\dots-(s+u+v)_{d+1}]\cdots[-u_{d+1}]$
Also $u_{r}v_{k}=0$ if $k<r$. Then, for each term, index $r$, for all $k<r$ remove $v_{k}$ from the $k$-th factor (before each $[-u_{k}]$).
$\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))=\dots=$
$[-u_{1}]\cdot[1-s_{1}-(s+u+v)_{2}-\dots-(s+u+v)_{d+1}]\cdots[1-s_{d}-(s+u+v)_{d+1}]$
$+[1-s_{1}-(s+u+v)_{2}-\dots-(s+u+v)_{d+1}]\cdot[-u_{2}]\cdots[1-s_{d}-(s+u+v)_{d+1}]$
+$\cdots$+ ⋯
$+[1-s_{1}-\dots-(s+u+v)_{d+1}]\cdot[1-s_{2}-\dots-(s+u+v)_{d+1}]\cdots[-u_{d+1}]$
Factors before and after $u_{r}$ are identical between terms. Factorize:
$\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))=\dots=$
$[u_{1}+u_{2}+\dots+u_{d+1}]$
$\quad\cdot[1-s_{1}-(s+u+v)_{2}-(s+u+v)_{3}-\dots-(s+u+v)_{d+1}]$
$\quad\cdot[1-s_{2}-(s+u+v)_{3}-\dots-(s+u+v)_{d+1}]$
[…]
$\quad\cdot[1-s_{d}-(s+u+v)_{d+1}]$
$s_{d+1}=0$, since $x_{E}$ is the only flat of rank $d+1$, and $i\in E$. That is, $E\notin\mathcal{S}$.
$\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))=\dots=$
$-[u_{1}+u_{2}+\dots+u_{d+1}]$
$\quad\cdot[1-s_{1}-(s+u+v)_{2}-(s+u+v)_{3}-\dots-(u+v)_{d+1}]$
$\quad\cdot[1-s_{2}-(s+u+v)_{3}-\dots-(u+v)_{d+1}]$
[…]
$\quad\cdot[1-s_{d}-(u+v)_{d+1}]$
$-[u_{1}+u_{2}+\dots+u_{d+1}]$
$\quad\cdot[1-(s_{1}+u_{2}+v_{2})-(s_{2}+u_{3}+v_{3})-\dots-(s_{d}+u_{d+1}+v_{d+1})]$
$\quad\cdot[1-(s_{2}+u_{3}+v_{3})-\dots-(s_{d}+u_{d+1}+v_{d+1})]$
[…]
$\quad\cdot[1-(s_{d}+u_{d+1}+v_{d+1})]$
$=-\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/{i})))$
As desired, we have
$\operatorname{stair}(M)=\bar{\delta}^{*}\operatorname{stair}(M\setminus i)-\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/{i})))$
@@ 4.2 Piecewise polynomials and cocycles
There is an equivalent formulation of $A^{*}(\Sigma)$ in terms of so-called Courant functions on $\Sigma$. This approach defines $A^{*}(\Sigma)$ as a quotient of continuous piecewise polynomial functions on $\Sigma$ by linear functions from $N$ to $\mathbb{R}$, see Section 4 of [AHK18] for more details.
This other formulation leads to the theory of tropical cocycles on $\Sigma$, introduced by François in [Fra13] and elaborated upon by Gross and Shokrieh in [GS19], in analogy to Allermann and Rau’s theory of cycles on $\Sigma$ from [AR09]. In particular, $Z_{*}(\Sigma)$, the tropical cycles on $\Sigma$, consists of weighted fans on subdivisions of $\Sigma$, while $C^{*}(\Sigma)$, the tropical cocycles on $\Sigma$, are continuous functions that are piecewise polynomial in a subdivision of $\Sigma$. In [GS19], the authors prove that, when given the corresponding intersection products, these $C^{*}(\Sigma)$-algebras are isomorphic. In particular, given a morphism of fans $f:\Sigma_{1}\to\Sigma_{2}$, a cocycle $h\in C^{*}(\Sigma_{2})$ and a cycle $X\in Z_{*}(\Sigma_{2})$, $f^{*}(h\cdot X)=f^{*}(h)\cdot f^{*}(X)$, where $f^{*}:Z_{*}(\Sigma_{2})\to Z_{*}(\Sigma_{1})$ is defined in Corollary 4.7 in [GS19], which agrees with Definition 2.16 in [Sha13] in the case where $f$ is a tropical modification.
@@
In the context of tautological classes of matroids:
Remark 4.1.
For an arbitrary matroid $M$, the class $\sum_{i=0}^{\operatorname{rk}(M)-1}c_{\operatorname{rk}(M)-1-k}(\mathcal{S_{M}})$ equals the staircase of $M$. In particular, the Chern roots of $S_{M}$ are
$\alpha_{r}=\sum_{\begin{subarray}{c}F\in\mathcal{L}(M)\\ \operatorname{rk}(F)\geq r\end{subarray}}x_{F}.$
Note that this is different, yet implied by Appendix III in [BEST23], as the use of modular filters is not needed and the computation happens in $A^{*}(M)$.
@@
The formula of Fife and Rincón holds for all matroids.

### v3

@@ 4 Relation to the CSM cycle of a matroid
Given a loopless matroid $M$, define the class $\operatorname{stair}(M)$ in its Chow ring $A^{*}(M)$ by the formula
@@
First we show how this class respects behaves with respect to certain deletion and contraction operations on a matroid.
@@
Let $M$ be a matroid on ground set $E$ with rank function $\operatorname{rk}_{M}$. For any $i\in E$, we can construct the matroid $M\setminus i$, called the deletion of $M$ by the element $i$. This is the matroid on the set $E\setminus i$ with rank function $\operatorname{rk}_{M\setminus i}:2^{E\setminus i}\to\mathbb{Z}$ given by:
@@
For any $i\in E$, we can construct the matroid $M/i$, called the contraction of $M$ by the element $i$. This is the matroid on the set $E\setminus i$ with rank function $\operatorname{rk}_{M\setminus i}:2^{E\setminus i}\to\mathbb{Z}$ given by:
@@ 4.1 Contraction-deletion of $\operatorname{stair}$
For a given element $i\in E$ which is not a coloop, we obtain a projection $\delta:B(M)\to B(M\setminus i)$. By Proposition 2.22 in [Sha13], $\delta$ is a tropical modification along some rational function $\phi$; therefore, $\delta$ induces a pullback $\delta^{*}:Z_{*}(B(M\setminus i)\to Z_{*}(B(M))$ and since $div_{B\setminus i}(\phi)=B(M/i)$, we also obtain a pullback $Z_{*}(B(M/i))\to Z_{*}(B(M\setminus i))$, where $Z_{*}(\Sigma)$ is the group of cycles on the fan $\Sigma$, see Definition 5.12 in [AR09]. These have corresponding ring/group homomorphisms in the context of the Chow rings, see Section 4.2 for details.
@@
To distinguish the generators of the Chow rings of $M,M\setminus i$ and $M/i$, we label them by
@@
$A^{*}(M/i)$ $=\mathbb{Z}[z_{F}\mid F\in\mathcal{L}(M/i)]/(I_{M/i}+J_{M/i}).$
Define the deletion pullback $\bar{\delta}^{*}:A^{*}(M\setminus i)\rightarrow A^{*}(M)$ as the ring homomorphism induced by
@@
where $\iota(z_{F})=y_{F}$, i.e. $\iota$ replaces the variables $z$ with $y$. This is an embedding $A^{*}(M/i)$ as the subgroup $y_{i}A^{*}(M\setminus i)$. By composition we also obtain a group homomorphism $\bar{\delta}^{*}(y_{i}\cdot\iota(\_)):A^{*}(M/i)\rightarrow A^{*}(M)$.
Using these definitions, the staircase class of a matroid satisfies the contraction-deletion formula stated below. For the proof of Theorem 4.1, see Section 5.
@@
For all non-coloop elements $i\in E$,
$\operatorname{stair}(M)=\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))-\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/i))).$
@@ 4.2 Piecewise polynomials and cocycles
There is an equivalent formulation of $A^{*}(\Sigma)$ in terms of so-called Courant functions on $\Sigma$. This approach defines $A^{*}(\Sigma)$ as a quotient of continuous piecewise polynomial functions with integer coefficients on $\Sigma$ by linear functions from $N$ to $\mathbb{R}$, see Section 4 of [AHK18] for more details.
This other formulation leads to the theory of tropical cocycles on $\Sigma$, introduced by François in [Fra13] and elaborated upon by Gross and Shokrieh in [GS19], in analogy to Allermann and Rau’s theory of cycles on $\Sigma$ from [AR09]. In particular, $Z_{*}(\Sigma)$, the tropical cycles on $\Sigma$, consists of weighted fans on subdivisions of $\Sigma$, while $C^{*}(\Sigma)$, the tropical cocycles on $\Sigma$, are continuous functions that are piecewise polynomial in a subdivision of $\Sigma$. In [GS19], the authors prove that, when given the corresponding intersection products, these $C^{*}(\Sigma)$-algebras are isomorphic. In particular, given a morphism of fans $f:\Sigma_{1}\to\Sigma_{2}$, a cocycle $\alpha\in C^{*}(\Sigma_{2})$ and a cycle $X\in Z_{*}(\Sigma_{2})$, $f^{*}(\alpha\cdot X)=f^{*}(\alpha)\cdot f^{*}(X)$, where $f^{*}:Z_{*}(\Sigma_{2})\to Z_{*}(\Sigma_{1})$ is defined in Corollary 4.7 in [GS19], which agrees with Definition 2.16 in [Sha13] in the case where $f$ is a tropical modification.
@@
The formula of Fife and Rincón holds for all matroids and the proofs of Corollary 7.0.1 and Proposition 7.0.4 in [Man22] are valid.
4.3 Connection to tautological classes of matroids
For an arbitrary set $E$, with $|E|=n$, the permutohedral variety $X_{E}$ is the (maximal) wonderful model of the hyperplane arrangement consisting of the coordinate hyperplanes in $\mathbb{P}^{n-1}$. The matroid associated to this hyperplane arrangement is the uniform matroid $U_{n,n}$ on $E$, for which all subsets are flats. Hence, its Chow ring $A^{*}(X_{E})$ is generated by variables $z_{F}$ for all subsets $\emptyset\subsetneq F\subset E$
For an arbitrary matroid $M$ on the same set $E$, the authors of [BEST23] define a tautological subbundle $\mathcal{S}_{M}$ and quotient bundle $\mathcal{Q}_{M}$ over $X_{E}$, with Chern classes valued in $A^{*}(X_{E})$.
If we denote by $\iota:A^{*}(M)\rightarrow A^{*}(X_{E})$ the map $\iota(x_{F})=z_{F}$, Theorem 7.6 in [BEST23] says that there is a canonical embedding $A^{*}(M)\rightarrow A^{*}(X_{E})$ as the ideal generated by $c_{|E|-\operatorname{rk}(M)}(\mathcal{Q}_{M})$, given by $c_{|E|-\operatorname{rk}(M)}(\mathcal{Q}_{M})\cdot\iota(\_)$.
Under this canonical identification, in Theorem 8.4 in [BEST23], the authors prove that the Poincaré dual of the CSM class is given by $c(\mathcal{S}_{M})\cdot c_{|E|-\operatorname{rk}(M)}(\mathcal{Q}_{M})$ and in Appendix III.1, they give the Chern root decomposition for $c(\mathcal{S}_{M})$, which yields the same staircase form $\iota(\operatorname{stair}(M))$ after removing the terms which are annihilated by $c_{|E|-\operatorname{rk}(M)}(\mathcal{Q}_{M})$.
Remark 4.1.
The map $c(\mathcal{S}_{M})\cdot\_:A^{*}(X_{E})\to A^{*}(X_{E})$ equals the map $\iota(\operatorname{stair}(M))\cdot\_:A^{*}(X_{E})\to A^{*}(X_{E})$, when restricted to the submodule $c_{|E|-\operatorname{rk}(M)}(Q_{M})A^{*}(X_{E})$.
5 Proof of Theorem 4.1
Proof.
We partition the nonempty flats of $M$ according to the flowchart:
[…]
Denote also $\mathcal{U}\setminus i:=\{F\setminus\{i\}|F\in\mathcal{U}\}$ and $\mathcal{V}\setminus i:=\{F\setminus\{i\}|F\in\mathcal{V}\}$.
Note the following reasons for this partition:
  • $\mathcal{U}\cup\mathcal{V}$ consists of all flats containing the deleted element $i$. Among these, the ones in $\mathcal{U}$ are the flats that decrease in rank when $i$ is deleted and the ones in $\mathcal{V}$ are the flats that remain the same rank when $i$ is deleted.
  • $\mathcal{U}\setminus i=\mathcal{S}\cup\{\emptyset\}$.
  • If $F\in\mathcal{T}$, then $\bar{\delta}^{*}(y_{F})=x_{F}$.
  • If $F\in\mathcal{S}$, then $\bar{\delta}^{*}(y_{F})=x_{F}+x_{F\cup\{i\}}$.
  • If $F\in\mathcal{V}\setminus i$, then $\bar{\delta}^{*}(y_{F})=x_{F\cup i}$.
Under the partition, the resulting quotient lattice structure is
[…]
For ranks $r=1,\dots,d+1$, let
One of $\mathcal{T},\mathcal{S},\mathcal{U}$ and $\mathcal{V}$ may be empty in some rank $r$, in which case the sum is zero. For short hand, we also group indices with the same rank $(t+s)_{r}:=t_{r}+s_{r}$, etc. In particular, note that
$\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{S}\\ \operatorname{rk}(F)=j\end{subarray}}y_{F}\right)$ $=s_{j}+u_{j+1},$ $\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{T}\\ \operatorname{rk}(F)=j\end{subarray}}y_{F}\right)$ $=t_{j}$
$\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{V}\setminus i\\ \operatorname{rk}(F)=j\end{subarray}}y_{F}\right)$ $=v_{j}$
Using this notation,
$\operatorname{stair}(M)$ $=\prod_{r=1}^{d+1}\left(1-\sum_{j=r}^{d+1}(s+t+u+v)_{j}\right)$
and
$\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))$ $=\prod_{r=1}^{d+1}\left(1-\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{L}(M\setminus i)\\ \operatorname{rk}(F)\geq r\end{subarray}}y_{F}\right)\right)$
$=\prod_{r=1}^{d+1}\left(1-\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{S}\\ \operatorname{rk}(F)\geq r\end{subarray}}y_{F}+\sum_{\begin{subarray}{c}F\in\mathcal{T}\\ \operatorname{rk}(F)\geq r\end{subarray}}y_{F}+\sum_{\begin{subarray}{c}F\in\mathcal{V}\setminus i\\ \operatorname{rk}(F)\geq r\end{subarray}}y_{F}\right)\right)$
$=\prod_{r=1}^{d+1}\left(1-\sum_{j\geq r}(s_{j}+u_{j+1})+\sum_{j\geq r}t_{j}+\sum_{j\geq r}v_{j}\right)$
$=\prod_{r=1}^{d+1}\left(1-(s+t+v)_{r}-\sum_{j=r+1}^{d+1}(s+t+u+v)_{j}\right)$
Note that $u_{r}$ is omitted from the $r$-th factor. We compute the difference $\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))$ and show it is equal to the additive inverse of the pullback to $A^{*}(M)$ of the staircase of $M/i$; concretely, denoting the rank function of $M$ by $\operatorname{rk}$, the rank function of $M/i$ by $\operatorname{rk}^{\prime}$ and noting that for $F\in\mathcal{V}\setminus i$, $\operatorname{rk}(F)=\operatorname{rk}^{\prime}(F)+1$,
$\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/i)))$ $=\bar{\delta}^{*}(y_{i})\cdot\bar{\delta}^{*}\iota(\operatorname{stair}(M/i)))$
$=\left(\sum_{r=1}^{d+1}u_{i}\right)\bar{\delta}^{*}\iota\left(\prod_{r=1}^{d}\left(1-\sum_{\begin{subarray}{c}F\in\mathcal{L}(M/i)\\ \operatorname{rk}^{\prime}(F)\geq r\end{subarray}}z_{F}\right)\right)$
$=\left(\sum_{r=1}^{d+1}u_{i}\right)\prod_{r=1}^{d}\left(1-\bar{\delta}^{*}\iota\left(\sum_{\begin{subarray}{c}F\in\mathcal{S}\\ \operatorname{rk}(F)\geq r\end{subarray}}z_{F}\right)-\bar{\delta}^{*}\iota\left(\sum_{\begin{subarray}{c}F\in\mathcal{V}\setminus i\\ \operatorname{rk}(F)\geq r+1\end{subarray}}z_{F}\right)\right)$
$=\left(\sum_{r=1}^{d+1}u_{i}\right)\prod_{r=1}^{d}\left(1-\bar{\delta}^{*}\iota\left(\sum_{\begin{subarray}{c}F\in\mathcal{S}\\ \operatorname{rk}(F)\geq r\end{subarray}}z_{F}\right)-\bar{\delta}^{*}\iota\left(\sum_{\begin{subarray}{c}F\in\mathcal{V}\setminus i\\ \operatorname{rk}(F)\geq r+1\end{subarray}}z_{F}\right)\right)$
$=\left(\sum_{r=1}^{d+1}u_{i}\right)\prod_{r=1}^{d}\left(1-\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{S}\\ \operatorname{rk}(F)\geq r\end{subarray}}y_{F}\right)-\bar{\delta}^{*}\left(\sum_{\begin{subarray}{c}F\in\mathcal{V}\setminus i\\ \operatorname{rk}(F)\geq r+1\end{subarray}}y_{F}\right)\right)$
$=\left(\sum_{r=1}^{d+1}u_{i}\right)\prod_{r=1}^{d}\left(1-\sum_{j\geq r}(s_{j}+u_{j+1})+\sum_{j\geq r}v_{j}\right)$
$=\left(\sum_{r=1}^{d+1}u_{i}\right)\prod_{r=1}^{d}\left(1-s_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)$
The difference is:
$-\sum_{k=1}^{d+1}\prod_{r=1}^{k-1}\left(1-(t+s+v)_{r}-\sum_{j=r+1}^{d+1}(t+s+u+v)_{j}\right)u_{k}\prod_{r=k+1}^{d+1}\left(1-\sum_{j=r}^{d+1}(t+s+u+v)_{j}\right)$
Note in the $k$-th term is like $\operatorname{stair}(M)$ except the $k$-th factor is replaced by $-u_{k}$ and for each $r<k$, $u_{k}$ is omitted from the $r$-th factors. Since all pairs of flats in $\mathcal{T}$ and $\mathcal{U}$ are not comparable, $t_{j}u_{l}=0$ for all $j$ and $l$. Since each term has some $u_{k}$ as a factor, all ‘$t$’s may be removed.
$-\sum_{k=1}^{d+1}\prod_{r=1}^{k-1}\left(1-(s+v)_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)u_{k}\prod_{r=k+1}^{d+1}\left(1-\sum_{j=r}^{d+1}(s+u+v)_{j}\right)$
Also $u_{k}s_{l}=0$ if $k\leq l$. Thus:
$-\sum_{k=1}^{d+1}\prod_{r=1}^{k-1}\left(1-(s+v)_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)u_{k}\prod_{r=k+1}^{d+1}\left(1-s_{r-1}-\sum_{j=r}^{d+1}(s+u+v)_{j}\right)$
Also $u_{k}v_{l}=0$ if $l<k$. Then, for each term, index $k$, for all $l<k$ remove $v_{l}$ from the first $l$-th factors (before each $-u_{k}$).
$\operatorname{stair}(M)-\bar{\delta}^{*}(\operatorname{stair}(M\setminus i))$
$=$ $-\sum_{k=1}^{d+1}\prod_{r=1}^{k-1}\left(1-s_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)u_{k}\prod_{r=k+1}^{d+1}\left(1-s_{r-1}-\sum_{j=r}^{d+1}(s+u+v)_{j}\right)$
$=$ $-\sum_{k=1}^{d+1}u_{k}\prod_{r=1}^{k-1}\left(1-s_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)\prod_{r=k}^{d}\left(1-s_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)$
$=$ $-\left(\sum_{k=1}^{d+1}u_{k}\right)\prod_{r=1}^{d}\left(1-s_{r}-\sum_{j=r+1}^{d+1}(s+u+v)_{j}\right)$
$=$ $-\bar{\delta}^{*}(y_{i}\cdot\iota(\operatorname{stair}(M/i))),$
which proves the claim. ∎